OpenStudy (anonymous):

Solve A=2(Π)r2+2(Π)rh for positive r." Hint: Use Quadratic Formula. Help? Medal will be given.

OpenStudy (kropot72):

Do you know the formula for solving any quadratic equation?

OpenStudy (jdoe0001):

\(\bf A=2\pi r^2+2\pi rh\implies A=2\pi (r^2+rh)\implies 0=2\pi (r^2+rh) \\ \quad \\ 0=r^2+rh\implies 0={\color{blue}{ 1}}r^2+{\color{red}{ h}}r{\color{green}{ +0}} \\ \quad \\ \textit{quadratic formula}\\ \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

Okay, I understand that part. But once I solved it and got \[x=(-h+h)/2\] Is that what r equals?? And in that case, the numerator always cancels to 0, correct? So r=0??

OpenStudy (accessdenied):

I think that is making an implicit assumption that A = 0, which is not generally true. Instead, we should treat A as another constant. \( A = 2 \pi r^2 + 2 \pi r h \) To get the general form of a quadratic, we subtract both sides by A. \( 0 = \color{red}{2 \pi } \ r^2 + \color{red}{2 \pi h} \ r \color{red}{- A} \) This is now a quadratic equation in r. Try not to be intimidated by the level of abstraction here: just work it down like any quadratic using the formula you already used, using a = \(2\pi\), b = \(2 \pi h\), and c = \(-A\).

OpenStudy (triciaal):

@AccessDenied how would you simplify =/- rt of [(2*pi*h)^2 + 8*pi*A} over 2*pi